# How does online shopping affect offline price sensitivity?

### v1

@@ Abstract
The rapid rise of e-commerce has transformed consumer behavior, prompting questions about how online adoption influences offline shopping. We examine whether consumers who adopt online shopping with a retailer become more price-sensitive in their subsequent offline purchases with that retailer. Using transaction-level data from a large Brazilian pet supplies retailer operating both online and offline, we compare “adopters”—customers who began shopping online after a period of offline-only purchasing—with “non-adopters” who remained offline-only. We estimate a discrete choice logit model with individual-level heterogeneity, using a novel algorithm to handle high-dimensional fixed effects and address price endogeneity. We then apply a staggered difference-in-differences approach to the estimated price elasticities and obtain the Average Treatment Effect on the Treated (ATT). We find that offline price sensitivity increases significantly after online adoption in three out of four product categories, particularly in items with low switching costs, such as pet hygiene. Counterfactual pricing simulations show that incorporating these behavioral spillovers into pricing strategies can increase firm profits by up to 4.1%. These results underscore the importance of recognizing cross-channel effects in consumer behavior and contribute to the literature on pricing and multichannel retailing by identifying online adoption as a key driver of offline price sensitivity.
Keywords: Multichannel Retailing, Online Shopping, Price Elasticity, Pricing
@@ 1 Introduction
Over the past two decades, online shopping has grown rapidly, driven by the widespread adoption of computers and smartphones. U.S. e-commerce sales more than doubled from \$571 billion in 2019 to approximately \$1.19 trillion in 2024, and the share of e-commerce in total US retail sales has grown from 15.4% to 22.7% over the same period (Haleem 2025). Consumers are increasingly adapting to the convenience of online shopping and allocating a large portion of their spending to online channels (McKinsey & Company 2023). This shift has prompted traditional brick-and-mortar retailers such as Walmart, Target, and Costco to expand into e-commerce channels and adopt multichannel strategies.
@@
Given the rise of online shopping, a natural question that arises for retailers is whether online shopping influences how consumers shop offline, and if so, how? In this paper, we focus on an important aspect of consumer behavior – their price sensitivity. Specifically, we seek to answer two key research questions – (1) Does adopting online shopping affect consumers’ offline price sensitivity? If so, does offline price sensitivity increase or decrease post-online adoption? (2) What are the implications of such cross-channel spillovers on retailers’ offline pricing strategies and profits? Given the migration of consumers from offline stores to multi-channel shopping, quantifying the existence and impact of cross-channel externalities (in particular, the effect of online shopping on offline shopping behavior) is important, both for improving our understanding of consumer behavior as well as for optimizing firm strategies in multi-channel settings.
@@
While there is a rich literature studying the determinants of price sensitivity, this literature largely focuses on determinants of price elasticities within a channel (see §$\S$2 for a detailed discussion). By contrast, our focus is on the externalities or spillover across channels. From a conceptual perspective, there are reasons why we may expect offline price sensitivity to increase post-online adoption as well as reasons that suggest the opposite. On the one hand, online shopping may make consumers less price sensitive and more brand loyal because online recommendation systems promote easy repeat purchases and lead to habit formation by highlighting prior purchases, e.g., “Buy again” or “Frequently purchased items” (Danaher et al. 2003; Chu et al. 2008). This increase in brand loyalty in online settings could spill over to offline shopping occasions and may make consumers more brand loyal and less price-sensitive offline too (after adopting online shopping). On the other hand, online shopping makes it easy for consumers to engage in price comparisons (due to features like sorting by price and lower search costs of browsing through many options on an app or website). This habit of checking and comparing prices online can potentially carry over to offline settings, and consumers become more aware of price differences in physical stores as well. Another potential reason that may make consumers more price-sensitive after adopting online shopping is that, in many contexts, prices are lower online than offline (Brynjolfsson and Smith 2000; Cooper 2006). Being exposed to lower prices for the same products online can make in-store prices seem inflated by comparison, prompting greater scrutiny of the prices of different brands. Finally, recent research suggests that when consumers are exposed to a wider variety of prices (which is likely to occur when they adopt online shopping), they become more price-sensitive (Aparicio et al. 2024). Therefore, the question of whether the adoption of online shopping affects offline price sensitivity, and if so, in which direction, is an empirical one.
@@
We address this question using data from a large pet supplies retailer in Brazil that sells through both online and offline channels. The dataset includes detailed transaction-level information for all customers across both channels from January 1, 2019, to June 30, 2021. In our data, we observe some consumers who were initially offline-only and made their first online purchase with the retailer, while others remained offline-only throughout. This allows us to track the offline purchases of adopters, both pre- and post-adoption, and also all purchases of those consumers who remained offline only. We focus on four major product categories: dog food, cat food, dog hygiene, and cat hygiene, which together account for more than 50% of the retailer’s total revenue. These categories are also relatively frequently purchased and are characterized by a good amount of price variation both across brands and across time, which allows us to measure individual-level price elasticities at the monthly level.
@@
Our empirical strategy consists of two steps. In the first stage, we specify and estimate a discrete choice logit model for each category, controlling for unobserved customer heterogeneity in brand preferences. In our model, we allow price elasticity to be a function of the extent to which a given consumer adopts online shopping at a point in time, measured by the cumulative fraction of their online spending with the retailer up to that point in time. This modeling approach allows us to measure the effect of interest and examine how the price elasticity of consumers who have adopted online shopping diverges over time from those who have not. We encounter two challenges in estimating this demand model. First, maximum likelihood estimation of a non-linear model, such as the logit model, is quite challenging in the presence of high-dimensional fixed effects (customer-brand fixed effects in our case). To accommodate these high-dimensional fixed effects, we build on the Minorization-Maximization algorithm proposed by Chen et al. (2021). We extend the Chen et al. (2021) algorithm (that uses a multinomial logit model with no alternative specific characteristics) to a more general setting where covariates vary by individual, by alternative, and also over time. We then apply this algorithm to our setting for estimation. Second, we must address potential endogeneity in prices using our observational data (Nevo 2001). To that end, we use costs as an instrument (Berry 1994) in conjunction with a control function approach (Villas-Boas and Winer 1999; Petrin and Train 2010). We then use the estimates from this model to compute customer-month-level price elasticities.
@@
In the second stage, we use the estimated customer-month-level price elasticities of adopters and non-adopters from the first stage to estimate our ATT of interest, which is the effect of adopting online shopping on customer price elasticities. Again, our observational data pose some challenges for the identification of the ATT. First, consumers in our sample are either treated or untreated in each time period, and we do not observe their counterfactual outcomes in the unobserved condition (i.e., their behaviors if they had not adopted online shopping). Moreover, the staggered timing of online adoption among different consumers requires careful computation and aggregation of treatment effects (Goodman-Bacon 2021; Callaway and Sant’Anna 2021; Baker et al. 2022). To deal with this, we employ a staggered difference-in-differences (DID) framework, as proposed by Callaway and Sant’Anna (2021), utilizing the variation in staggered adoption timing of online shopping. We identify different cohorts of consumers who adopt online shopping (i.e., adopters) in each month of our data and compare changes in their behavior to those of a control group of customers who shop exclusively offline throughout the observation period (i.e., non-adopters).
@@
Our main finding is that consumers become more price sensitive in their offline shopping trips after adopting online shopping for three out of four categories. The relative increase in offline price elasticity, defined as the ATT relative to the pre-adoption average price elasticity of the treated group is heterogeneous across categories. We find a large effect in the dog and cat hygiene categories (where price elasticity increases by 72.7% and 34.7%, respectively). We also observe a significant but smaller increase in the dog food category (8.4%) and no effect for cat food. The heterogeneity across categories can likely be explained by differences in switching costs to different brands. While pet owners face little risk in trying a new lower-priced hygienic mat for their pet, changing the brand of food can be much trickier due to the pets’ stronger preferences for shape, smell, and texture in food, particularly for cats (Nichelason 2024). We also provide a series of tests and alternative specifications that establish the robustness of our results and provide additional empirical support for the generalizability of our findings.
Our findings have significant managerial implications for firms’ pricing strategies. Based on our empirical results, we consider counterfactual pricing scenarios and document the value of correctly accounting for the online adoption effect on offline prices. We first consider a uniform pricing strategy, where all customers pay the same price, and the firm correctly accounts for the online adoption effect when optimizing prices. We demonstrate that this approach can increase the firm’s offline profit (summed across the three categories where there is a significant online adoption effect) by 1.17%, compared to a scenario where the manager does not account for the online adoption effect in their demand model and attempts to set optimal offline prices. Next, we perform a similar comparison under a personalized pricing scenario, where the firm can personalize prices (similar to Rossi et al. (1996); Smith et al. (2023); Jain et al. (2024)). We show that incorporating the customer’s online spend in the behavioral features in the demand model and optimizing prices based on this demand model increase profits by 4.1% (compared to a setting where the firm does not include the online adoption effect in the demand model but still personalizes prices). Together, these counterfactual results suggest correctly accounting for online spend in offline demand models can help retailers improve their pricing and personalization strategies.
@@
In sum, our paper makes two main contributions to the literature. First, from a substantive perspective, we contribute to the literature on price sensitivity and multichannel retailing by identifying a new determinant of offline price sensitivity: the adoption of online shopping. We demonstrate that consumers become more price-sensitive in the offline channel after adopting online shopping. This finding is important and distinct from previous research, which has mainly focused on cross-sectional or within-household comparisons of price sensitivity between online and offline settings (Degeratu et al. 2000; Chu et al. 2008, 2010). Second, from a managerial perspective, our counterfactual simulations suggest that incorporating the online adoption effects into price elasticity measurements and subsequent price optimization can significantly improve retailer profits. Overall, these findings highlight the importance of correctly accounting for cross-channel spillovers in retailers’ marketing decisions.
@@
The rest of the paper is organized as follows. In §2, we discuss the related literature, and in §3, we discuss the setting and data. In §4, we document descriptive evidence, and in §5, we specify the firm’s problem and provide an overview of the empirical strategy. §6 presents the demand model, and §7 presents the ATT estimates of online adoption on offline price sensitivity. §8 discusses the generalizability and robustness checks of our main results. In §9, we perform counterfactual simulations and provide managerial implications for retailers’ offline pricing and profitability. Finally, in §10, we summarize our findings and discuss some directions for future research.
2 Literature
@@
Our study relates to multiple streams of literature on the determinants of price elasticity, optimal pricing, and multi-channel retailing in marketing and economics.
@@
First, our work relates to the large literature in marketing and economics that has focused on documenting the drivers of consumers’ price sensitivity. Prior research has identified various determinants of price sensitivity, particularly in grocery settings. These include (1) macroeconomic conditions, such as GDP growth and inflation rates (Gordon et al. 2013; Tellis 1988; Bijmolt et al. 2005); (2) marketing interventions, such as advertising (Kanetkar et al. 1992; Kaul and Wittink 1995; Kalra and Goodstein 1998; Jedidi et al. 1999) and promotions Mela et al. (1997); Jedidi et al. (1999); Kopalle et al. (1999); (3) brand strength (Guadagni and Little 1983; Kamakura and Russell 1989; Guadagni and Little 2008); (4) consumer loyalty (Danaher et al. 2003); (5) time between purchase and consumption (Joo et al. 2020). Early research has also examined the relationship between customer demographic characteristics – age, education, income, and household size – and price sensitivity using household-level scanner panel data (Rossi and Allenby 1993; Narasimhan 1984; Bawa and Shoemaker 1987, 1989). However, later studies have concluded these relationships are generally weak, suggesting that demographics do not explain a large portion of the variation in household-level price sensitivity (Hoch et al. 1995; Rossi et al. 1996).
@@
We consider the adoption of online shopping as a novel factor that may affect consumers’ offline price sensitivity. Given the rise of digitization and online shopping, this factor can be particularly important for multi-channel retailers (who are increasingly investing in online channels). Existing literature provides some insights into how consumer behavior differs between online and offline channels. Early studies focusing on cross-sectional comparisons find that consumers exhibit lower price sensitivity and less brand switching when shopping online compared to offline (Degeratu et al. 2000; Andrews and Currim 2004). Danaher et al. (2003) shows that high market share brands exhibit greater brand loyalty online, whereas low market share brands exhibit greater brand loyalty in offline stores. However, these studies are “cross-sectional” in nature, i.e., they compare the set of consumers shopping offline vs. those shopping online, who may be very different from each other. More recent research has shifted toward within-household analysis. For example, Chu et al. (2008) compare price sensitivity in online and offline grocery shopping within the same household and find that consumers are generally less price-sensitive online. Further, Chu et al. (2010) investigate the role of household characteristics in moderating price sensitivity across channels. However, none of this research looks into if and how the adoption of online shopping affects consumers’ offline price sensitivity, which is the focus of our paper. This is an important question since consumers in many traditional brick-and-mortar settings are increasingly adopting online shopping and becoming multi-channel shoppers. If the adoption of online shopping systematically changes consumers’ offline shopping behavior, then firms need to account for these changes in their pricing and promotion strategies.
Our research also relates to the literature on how price sensitivity estimates can affect firms’ strategic business decisions and policy interventions. A recent stream of work has looked at how we can use consumer-level demand models to design targeted pricing and promotion strategies to maximize profit (Smith et al. 2023; Dubé and Misra 2023; Jain et al. 2024). We build on this work and show that accurately accounting for consumers’ online adoption status can improve firms’ pricing decisions and profitability. More broadly, from a policy and research perspective, accurate measures of price sensitivity can help policy-makers and managers assess market power, competitive dynamics, and the welfare implications of price targeting (Kadiyali et al. 2000; Nevo 2001; Chintagunta et al. 2002; Dubé and Misra 2023; Beresteanu et al. 2024). While we do not focus on public policy and competition in this work, our paper suggests that the adoption of online shopping is an important factor affecting price sensitivity that policymakers and researchers should consider when making policy recommendations in multi-channel settings.
@@
Finally, our study contributes to the broader multichannel retailing literature; see Zhang et al. (2010) and Cui et al. (2021) for comprehensive overviews. One key stream of research highlights that online and offline channels can function as complements (Avery et al. 2012; Wang and Goldfarb 2017; Bell et al. 2018) or substitutes (Forman et al. 2009; Li 2021; Shriver and Bollinger 2022), depending on the context. Other studies have examined how digitization affects consumer shopping behavior across channels, focusing on spend patterns and purchase volume (Wang et al. 2015; Narang and Shankar 2019; Biswas et al. 2024), brand exploration (Pozzi 2012), and purchase variety (Chintala et al. 2023). Building on this body of work, our study provides new insights into multichannel strategy (Neslin 2022) by investigating changes in offline price sensitivity when consumers adopt online shopping.

### v2

@@ Abstract
The rapid rise of e-commerce has transformed consumer behavior, prompting questions about how online adoption influences offline shopping. We examine whether consumers who adopt a retailer’s online shopping channels become more price-sensitive in their subsequent offline purchases with that retailer. Using transaction-level data from a large Brazilian pet supplies retailer operating both online and offline, we compare “adopters”—customers who began shopping online after a period of offline-only purchasing—with “non-adopters” who remained offline-only. We estimate a discrete choice logit model with individual-level heterogeneity, based on an algorithm that can handle both high-dimensional fixed effects and price endogeneity. We then apply a staggered difference-in-differences approach to the estimated price elasticities and obtain the Average Treatment Effect on the Treated (ATT). We find that offline price sensitivity increases significantly after online adoption in three out of four product categories, particularly in items with low switching costs, such as pet hygiene. These results underscore the importance of recognizing cross-channel effects in consumer behavior and contribute to the literature on pricing and multichannel retailing by identifying online adoption as a key driver of offline price sensitivity.
Keywords: Digital Platforms, High-dimensional fixed effects, Multichannel Retailing, Online Shopping, Price Elasticity, Pricing
@@ 1 Introduction
Over the past two decades, online shopping has grown rapidly, driven by the widespread adoption of computers and smartphones. U.S. e-commerce sales more than doubled from \$571 billion in 2019 to approximately \$1.19 trillion in 2024, and the share of e-commerce in total US retail sales has grown from 15.4% to 22.7% over the same period (Haleem, 2025). Consumers are increasingly adapting to the convenience of online shopping and allocating a large portion of their spending to online channels (McKinsey & Company, 2023). This shift has prompted traditional brick-and-mortar retailers such as Walmart, Target, and Costco to increase investments into their e-commerce channels as a major driver of growth.
@@
Given the rise of online shopping and investments in e-commerce, a natural question that arises for retailers is whether a consumer’s adoption of a given retailer’s online channels influences her shopping behavior at that retailer’s physical stores, and if so, how? In this paper, we focus on one important aspect of consumer behavior – their price sensitivity. Specifically, we seek to answer the following research question: When an offline-only consumer starts using the online shopping channel(s) of a given retailer, does it affect her offline price sensitivity (i.e., price sensitivity at that retailer’s physical stores)? If so, does offline price sensitivity increase or decrease post-online adoption? There are two main reasons why answering this question is important from both researchers’ and managers’ perspectives. First, from a substantive perspective, understanding the determinants of price elasticity is important – doing so gives us insights into what factors make consumers more or less price sensitive, and how. Second, from a managerial perspective, quantifying the spillover effects of online shopping on offline price sensitivity can have significant implications for retailers’ pricing strategies and profits. Given the migration of consumers from offline stores to multi-channel shopping, quantifying the existence and impact of cross-channel externalities can help firms optimize and coordinate pricing and promotion across channels. More broadly, there is a growing interest among managers in understanding how interactions and spillovers between online and offline shopping channels shape consumer behavior, given the growth of omnichannel retailing across the globe (Zhang et al., 2019; Biswas et al., 2024).
@@
While there is a rich literature studying the determinants of price sensitivity, this literature largely focuses on determinants of price elasticities within a channel (see $\S$2 for a detailed discussion). In contrast, our focus is on the externalities or spillovers across channels. From a conceptual perspective, there are reasons why we may expect offline price sensitivity to increase post-online adoption as well as reasons that suggest the opposite. On the one hand, online shopping may make consumers less price sensitive and more brand loyal because online recommendation systems promote easy repeat purchases and lead to habit formation by highlighting prior purchases, e.g., “Buy again” or “Frequently purchased items” (Danaher et al., 2003). This increase in brand loyalty in a retailer’s online channels could spill over to offline shopping occasions and may make consumers more brand loyal and less price-sensitive in that retailer’s physical stores too (after adopting online shopping). On the other hand, online shopping makes it easy for consumers to engage in price comparisons (due to features like sorting by price and lower search costs of browsing through many options on an app or website). This habit of checking and comparing prices on a retailer’s online channels can potentially carry over to the retailer’s offline channels, and consumers may become more aware of price differences in the retailer’s physical stores as well. Another potential reason that may make consumers more price-sensitive after adopting online shopping is that, in many contexts, prices are lower online than offline (Brynjolfsson and Smith, 2000; Cooper, 2006). Being exposed to lower prices for the same products online can make in-store prices seem inflated by comparison, prompting greater scrutiny of the prices of different brands. Finally, recent research suggests that when consumers are exposed to a wider variety of prices (which is likely to occur when they adopt online shopping), they become more price-sensitive (Aparicio et al., 2024). Therefore, the question of whether the adoption of a retailer’s online channels affects consumers’ offline price sensitivities at that retailer’s physical stores, and if so, in which direction, is an empirical one.
@@
We address this question using data from a large pet supplies retailer in Brazil that sells through both online and offline channels. The dataset includes detailed transaction-level information for all customers across both channels from January 1, 2019, to June 30, 2021. In our data, we observe some consumers who were initially offline-only and then made their first online purchase with the retailer within our data window, while others remained offline-only throughout. This allows us to track the offline purchases of online shopping adopters, both pre- and post-adoption, and also all purchases of those consumers who remained offline only. We focus on four major product categories: dog food, cat food, dog hygiene, and cat hygiene, which together account for more than 50% of the retailer’s total revenue. These categories are also relatively frequently purchased and are characterized by a good amount of price variation both across brands and across time, which allows us to measure individual-level price elasticities at the monthly level.
@@
Our empirical strategy consists of two steps. In the first stage, we specify and estimate a discrete choice logit model for each category, controlling for unobserved customer heterogeneity in brand preferences. In our model, we allow consumers’ offline price sensitivity to be a function of the extent to which a given consumer adopts online shopping at a point in time, measured by the cumulative fraction of their online spending with the retailer up to that point in time. [^1] We also include an alternate specification with a binary online adoption variable, as a robustness check. This modeling approach allows us to measure the effect of interest and examine how the price elasticity of consumers who have adopted online shopping diverges over time from that of those who have not. We encounter two challenges in estimating this demand model. First, maximum likelihood estimation of a non-linear model, such as the logit model, is quite challenging in the presence of high-dimensional fixed effects (customer-brand fixed effects in our case). To accommodate these high-dimensional fixed effects, we build on the Minorization-Maximization algorithm proposed by Chen et al., (2022) and allow for covariates that vary by individual, by alternative, and also over time. We then apply this algorithm to our setting for estimation. To the best of our knowledge, this is the first empirical paper in marketing to efficiently estimate high-dimensional customer-brand fixed effects in a demand model using the MM algorithm. Second, we must address potential endogeneity in prices using our observational data (Nevo, 2001). To that end, we use costs as an instrument (Berry, 1994) in conjunction with a control function approach (Villas-Boas and Winer, 1999; Petrin and Train, 2010). We then use the estimates from this model to compute customer-month-level price elasticities.
@@
In the second stage, we use the estimated customer-month-level offline price elasticities of adopters and non-adopters from the first stage to estimate our ATT of interest, which is the effect of adopting online shopping on consumers’ offline price elasticities. Again, our observational data pose some challenges for the identification of the ATT. First, consumers in our sample are either treated or untreated in each time period, and we do not observe their counterfactual outcomes in the unobserved condition (i.e., their behaviors if they had not adopted online shopping). Moreover, the staggered timing of online adoption among different consumers requires careful computation and aggregation of treatment effects (Goodman-Bacon, 2021; Callaway and Sant’Anna, 2021; Baker et al., 2022). To deal with this, we employ a staggered difference-in-differences (DID) framework, as proposed by Callaway and Sant’Anna, (2021), utilizing the variation in staggered adoption timing of online shopping. We identify different cohorts of consumers who adopt online shopping (i.e., adopters) in each month of our data and compare changes in their behavior to those of a control group of customers who shop exclusively offline throughout the observation period (i.e., non-adopters).
@@
Our main finding is that consumers become more price sensitive in their offline shopping trips after adopting online shopping for three out of four categories. The relative increase in offline price elasticity, defined as the ATT relative to the pre-adoption average price elasticity of the treated group is heterogeneous across categories. We find a large effect in the dog and cat hygiene categories (where price elasticity increases by 72.7% and 34.7%, respectively). We also observe a significant but smaller increase in the dog food category (8.4%) and no effect for cat food. The heterogeneity across categories can likely be explained by differences in switching costs to different brands. While pet owners face little risk in trying a new lower-priced hygienic mat for their pet, changing the brand of food can be much trickier due to the pets’ stronger preferences for shape, smell, and texture in food, particularly for cats (Nichelason, 2024).
We also provide a series of tests and alternative specifications that establish the robustness of our results and provide additional empirical support for the generalizability of our findings. First, we show that using a binary online adoption variable in our demand model does not change our conclusions. Second, we also use Inverse Propensity Weighting (IPW) approaches to control for self-selection of the online adoption decision. Third, we consider an alternative estimator where later adopter groups serve as a control group for earlier adopters. Fourth, we consider an alternate definition of purchase occasions to also include online shopping visits. Our conclusions and substantive findings remain unchanged across all of these robustness checks. More broadly, we also present some evidence in favor of portability to other settings. For instance, our data period also includes COVID-19, an exogenous macro-environmental shock. Interestingly, we do not find differences in the adoption effect between customers who adopted online shopping during COVID-19 vs. those who adopted organically before that, indicating that the effect is not sensitive to the reasons/drivers of a consumer’s decision to adopt online shopping.
@@
Our paper contributes to the broad literature on online-offline retailing and pricing. From a substantive perspective, we contribute to the literature on price sensitivity and multichannel retailing by identifying a new determinant of offline price sensitivity: the adoption of online shopping. We demonstrate that consumers become more price-sensitive in a retailer’s physical stores after adopting that retailer’s online channels for shopping. This finding is important and distinct from previous research, which has mainly focused on cross-sectional or within-household comparisons of price sensitivity between online and offline settings. From a managerial perspective, our findings have implications for firms’ pricing strategies and profitability, emphasizing the need for multichannel retailers to account for the online adoption effect (i.e., heightened offline price sensitivity when consumers adopt a retailer’s online channels) in their pricing decisions.
@@
The rest of the paper is organized as follows. In §2, we discuss the related literature, and in §3, we discuss the setting and data. In §4, we document descriptive evidence, and in §5, we specify the firm’s problem and provide an overview of the empirical strategy. §6 presents the demand model, and §7 presents the ATT estimates of online adoption on offline price sensitivity. §8 discusses the generalizability and robustness checks of our main results. Finally, in §9, we summarize our findings and their implications, and also discuss avenues for future research.
2 Related Literature
@@
Our study relates to multiple streams of literature on the determinants of price elasticity, digital platforms, and multi-channel retailing.
@@
First, our work relates to the large literature in marketing and economics that has focused on documenting the drivers of consumers’ price sensitivity. Prior research has identified various determinants of price sensitivity, particularly in grocery settings. These include (1) macroeconomic conditions, such as GDP growth and inflation rates (Gordon et al., 2013; Tellis, 1988; Bijmolt et al., 2005); (2) marketing interventions, such as advertising (Kanetkar et al., 1992; Kaul and Wittink, 1995; Kalra and Goodstein, 1998; Jedidi et al., 1999), promotions Mela et al., (1997); Jedidi et al., (1999); Kopalle et al., (1999) and category expansion (Yu et al., 2021); (3) brand strength (Guadagni and Little, 1983; Kamakura and Russell, 1989; Guadagni and Little, 2008); (4) consumer loyalty (Danaher et al., 2003); (5) time between purchase and consumption (Joo et al., 2020). Early research has also examined the relationship between customer demographic characteristics – age, education, income, and household size – and price sensitivity using household-level scanner panel data (Rossi and Allenby, 1993; Narasimhan, 1984; Bawa and Shoemaker, 1987, 1989). However, later studies have concluded these relationships are generally weak, suggesting that demographics do not explain a large portion of the variation in household-level price sensitivity (Hoch et al., 1995; Rossi et al., 1996).
@@
We consider the adoption of a given retailer’s online channels as a novel factor that may affect consumers’ offline price sensitivity at that retailer’s physical stores. Given the rise of digitization and online shopping, this factor can be particularly important for multi-channel retailers (who are increasingly investing in online channels). Existing literature provides some insights into how consumer behavior differs between online and offline channels. Early studies focusing on cross-sectional comparisons find that consumers exhibit lower price sensitivity and less brand switching when shopping online compared to offline (Degeratu et al., 2000; Andrews and Currim, 2004). Danaher et al., (2003) shows that high market share brands exhibit greater brand loyalty online, whereas low market share brands exhibit greater brand loyalty in offline stores. However, these studies are “cross-sectional” in nature, i.e., they compare the set of consumers shopping offline vs. those shopping online, who may be very different from each other. More recent research has shifted toward within-household analysis. For example, Chu et al., (2008) compare price sensitivity in online and offline grocery shopping within the same household and find that consumers are generally less price-sensitive online. Further, Chu et al., (2010) investigate the role of household characteristics in moderating price sensitivity across channels. However, none of this research looks into if and how the adoption of online shopping affects consumers’ offline price sensitivity, which is the focus of our paper.
@@
More broadly, our study contributes to the broader multichannel retailing literature; see Cui et al., (2021) for a comprehensive overview. One key stream of research highlights that online and offline channels can function as complements (Avery et al., 2012; Wang and Goldfarb, 2017; Bell et al., 2018) or substitutes (Forman et al., 2009; Li, 2021; Shriver and Bollinger, 2022), depending on the context. Other studies have examined how digitization affects consumer shopping behavior across channels, focusing on spend patterns and purchase volume (Wang et al., 2015; Narang and Shankar, 2019; Biswas et al., 2024), brand exploration (Pozzi, 2012), and purchase variety (Chintala et al., 2023). Building on this body of work, our study provides new insights into multichannel strategy (Neslin, 2022) by investigating changes in offline price sensitivity when consumers adopt a retailer’s online channels for shopping.